OpenStudy (anonymous):

Melissa is canning tomato sauce. Currently she has the sauce filling a coontainer that has a volume of 9,000 cm3.How many cans with a height of 12 cm and a diameter of 10 cm will she need to hold all of her sauce?

OpenStudy (anonymous):

So here is the my solution. We know the can is a cylinder and the volume equation for a cynlinder is V=(pi)*(radius)^2*height. The diameter is 10 cm therefore the radius is 5cm. Pi is assumed to be 3.14. So using the above equation Volume = 3.14* (5)^2*12 thus V = (3.14)(25)12. V=942cm^3 Given the container volume is 9000cm^3 then 9000/942=9.6 cans.